OpenStudy (anonymous):

systems of equations: x/3+y/4=1 4x+3y=12 please explain so that I can learn what I am doing.

OpenStudy (mayankdevnani):

\[\large \bf \frac{x}{3}+\frac{y}{4}=1\] Take LCM, \[\large \bf \frac{4x+3y}{12}=1\]

OpenStudy (mayankdevnani):

so, \[\large \bf 4x+3y=12.................(1)\] \[\large \bf 4x+3y=12.....................(2)\]

OpenStudy (mayankdevnani):

now,both the equations are same,so what can you say about ? @Chubba

OpenStudy (anonymous):

that it would be infinite solutions because it will come out to be 0=0

OpenStudy (mayankdevnani):

yeah !!!

OpenStudy (mayankdevnani):

so any doubts ? @Chubba

OpenStudy (anonymous):

no thank you, i appreciate it. (my teacher keeps messing words up so confusing me) and now i will be able to take the test.

OpenStudy (mayankdevnani):

welcome :)